OpenStudy (anonymous):

Verify the expression: - tan^2x + sec^2x = 1

OpenStudy (abb0t):

HINT: \[\tan(x) = \frac{ \sin(x) }{ \cos(x) }, \sec(x) = \frac{ 1 }{ \cos(x) }\]

OpenStudy (anonymous):

(1/cosx^(2))-(sinx/sinx)^(2)?

OpenStudy (anonymous):

sorry i meant (1/cosx^(2))-(sinx/cosx)^(2)?

OpenStudy (abb0t):

Combine everything into one whole fraction. It might help you see where to go from there.

OpenStudy (anonymous):

(1-sinx^2)/(cosx^2)

OpenStudy (abb0t):

HINT #2: \[\sin^2(x)-1 = \cos^2(x)\]

OpenStudy (anonymous):

(1-sin^(2)x)/(sin^(2)x-1) ??

OpenStudy (abb0t):

You could do that too. you should see that the proof is now complete.

OpenStudy (abb0t):

Also: \[\frac{ \cos^2(x) }{ \cos^2(x) }\]

OpenStudy (anonymous):

How is 1-sin^(2)x and sin^(2)x-1 the same?

OpenStudy (anonymous):

how does 1-sin^(2)x also equal cos^(2)x?

OpenStudy (abb0t):

Yes, they are the same. So therefore, what do you get when you have two of the same?? AND Because: \[\sin^2(x)+\cos^2(x) = 1\] rearranging it gives you: \[\sin^2(x)-1 = \cos^2(x)\]all I did was rearrange it using basic algebra.

OpenStudy (anonymous):

OH! Okay so since tope and bottom both equal cos^(2)x, it's 1! Thank you so much!

OpenStudy (anonymous):

top*

OpenStudy (abb0t):

well, you changed the denomiator to match the top so yeah. but i was thinking of changing the numerator to the denominator. but either way works and gives you 1 = 1 hence, the proof holds true.

OpenStudy (anonymous):

Okay cool! thanks again!